# One-dimensional lattice random walks in a Gaussian random potential

### v2

@@ 6 The realization-dependent diffusion coefficient
In this Section we seek for the analogues of the result in eq. (71) for the random walks on discrete periodic chains with $N$ sites and also for different types of dynamics, as defined in our Models I, II and III. To this end, we will make use of Derrida’s general, though rather intricate, expression for the effective diffusion coefficient $D_{N}$, valid for an arbitrary set of transition rates [4]. In B we present details of the derivation of $D_{N}$ for the three models under study. Then, we will determine the moments of $D_{N}$, discuss their dependence on $\beta\sigma$ and on this basis will demonstrate that the realization-dependent diffusion coefficient is self-averaging in the limit $N\to\infty$, regardless of the dynamical scenario.
@@ 6.1 Model I: Random force-like model
In B we show that for such a model the Derrida’s formula [4] reduces to the following expression
@@ 6.2 Model II: Random walks with randomized stepping-times
In B we show that for this model the realization-dependent diffusion coefficient obeys the following symmetric form
@@ 6.3 Model III: Gaussian random trap model
We show in B that within this dynamical scenario the realization-dependent effective diffusion coefficient attains a particularly simple form:
@@
The auxiliary function $\Theta^{(III)}(\lambda)$ can be calculated very straightfowardly to give
@@
Figure 8: Numerically-evaluated normalized probablity density functions of the effective diffusion coefficients for the Models I (left panel), II (middle panel) and III (right panel) for different values of NN ranging from $N=16$ to $N=1024$ (see the insets). Vertical dashed lines indicate the limiting ($N\to\infty$) values of the disoder-averaged diffusion coefficients (see eqs. (79), (86) and (93) with $q=1$).
@@ 7 Conclusions
For all these random variables we have calculated analytically the leading (in the large-$N$ limit) asymptotic behavior of their moments determining the dependence on the strength of disorder and the temperature. We have shown that these moments exhibit an interesting super-Arrhenius dependence on the temperature, which is also quite different depending on the dynamical scenario. Comparing the averaged and the typical behavior we have demonstrated that the moments of the resistance, currents and the splitting probability are supported by atypical realizations of disorder.
@@
Further on, we have addressed a conceptually important question whether these transport properties exhibit self-averaging with respect to different realizations of disorder in the limit $N\to\infty$. Capitalizing on the derived exact expressions for the moments, we have studied the asymptotic behavior of their relative variances. We have shown that the probability currents through finite chains, the resistances of these chains and the splitting probabilities are not self-averaging in the limit $N\to\infty$. Interestingly enough, such a lack of self-averaging appears to be a boundary effect; namely, it stems from fluctuations of one (or a few) sites close to the boundary, while the contributions comming from the sites in the bulk of chains self-averages.
Concurrently, we have shown that the realization-dependent mean first-passage times to a target placed at distance $N$ away from the starting point are self-averaging in the limit $N\to\infty$. Clearly, this result just indicates the trend in the behavior of fluctuations that they become progressively less important the farther is the target away from the starting point. However, for most of practical applications $N$ is essentially finite and our numerical analysis shows that for such $N$ fluctuations are quite significant.
@@ Acknowledgments
The authors wish to thank E. Agliari, T. Akimoto, D. S. Dean and L. A. Pastur for helpful discussions. This project has been supported by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783—SMaC—Statistical Mechanics and Complexity University and Research) from Italian MUR (Ministry of University and Research).
@@ Appendix A Realization-dependent probability current
For the Model I the rates are defined in eqs. (LABEL:rates), which gives straightforwardly ‘
@@
from which we readily find the first expression in eq. (21).
@@
Appendix B Realization-dependent diffusion coefficient
@@
Derrida’s result
For completeness, we present below the Derrida’s general result for the realization-dependent $D_{N}$ for random walks evolving in a continuous time on a finite periodic chain with $N$ sites with arbitrary fixed transition rates $W_{x,x\pm 1}$. As shown by Derrida [4], the realization-dependent diffusion coefficient in such a setting has the following form
@@
Model I
For the Model I the transition rates are defined in eq. (LABEL:rates); that being,
$W_{x,x\pm 1}=W_{0}\frac{\phi_{x}}{\phi_{x\pm 1}}.$ (107)
Using this definition, we find
@@
where $H_{0,N}$ is given by
$H_{n,n+N}$ $\equiv\sum_{i=0}^{N-1}\phi_{n+i}\phi_{n+1+i}=H_{0,N}\,.$ (110)
Further on, from eq. (109) we have that
@@
Model II
In this dynamical scenario with randomized stepping-times, the transition rates $W_{x,x\pm 1}=p_{x,x\pm 1}/\delta t$, where $p_{x,x\pm 1}$ are the transition probabilities of an auxiliary random process that obey eqs. (11) and (13):
$p_{x,x\pm 1}=\frac{\phi_{x\mp 1}}{\phi_{x-1}+\phi_{x+1}}.$ (116)
Firstly, we notice that
@@
Model III
In the Gaussian random trap model the transition rates obey eq. (20), i.e.,
$W_{x,x\pm 1}=W_{0}\phi_{x}^{2}.$ (124)
It is straighforward to verify that the product

### v3

@@ 6 The realization-dependent diffusion coefficient
In this Section we seek for the analogues of the result in Eq. (68) for the random walks on discrete periodic chains with $N$ sites and also for different types of dynamics, as defined in our Models I, II and III. To this end, we will make use of Derrida’s general, though rather intricate, expression for the effective diffusion coefficient $D_{N}$, valid for an arbitrary set of transition rates [4]. In C we present details of the derivation of $D_{N}$ for the three models under study. Then, we will determine the moments of $D_{N}$, discuss their dependence on $\beta\sigma$ and on this basis will demonstrate that the realization-dependent diffusion coefficient is self-averaging in the limit $N\to\infty$, regardless of the dynamical scenario.
@@ 6.1 Model I: Random force-like model
In C we show that for such a model the Derrida’s formula [4] reduces to the following expression
@@ 6.2 Model II: Random walks with randomized stepping-times
In C we show that for this model the realization-dependent diffusion coefficient obeys the following symmetric form
@@ 6.3 Model III: Gaussian random trap model
We show in C that within this dynamical scenario the realization-dependent effective diffusion coefficient attains a particularly simple form:
@@
The auxiliary function $\Theta^{(III)}(\lambda)$ can be calculated very straightforwardly to give
@@
Figure 8: Numerically-evaluated normalized probability density functions of the effective diffusion coefficients for the Models I (left panel), II (middle panel) and III (right panel) for different values of NN ranging from $N=16$ to $N=1024$ (see the insets). Vertical dashed lines indicate the limiting ($N\to\infty$) values of the disorder-averaged diffusion coefficients (see Eqs. (76), (83) and (90) with $q=1$).
@@ 7 Conclusions
For all these random variables we have calculated analytically the leading (in the large-$N$ limit) asymptotic behavior of their moments determining the dependence on the strength of disorder and the temperature. We have shown that these moments exhibit an interesting super-Arrhenius dependence on the temperature, which is also quite different depending on the dynamical scenario. Comparing the averaged and the typical behavior we have demonstrated that the moments of the resistance and currents are supported by atypical realizations of disorder.
@@
Further on, we have addressed a conceptually important question whether these transport properties exhibit self-averaging with respect to different realizations of disorder in the limit $N\to\infty$. Capitalizing on the derived exact expressions for the moments, we have studied the asymptotic behavior of their relative variances. We have shown that the probability currents through finite chains and the resistances of these chains are not self-averaging in the limit $N\to\infty$. Interestingly enough, such a lack of self-averaging appears to be a boundary effect; namely, it stems from fluctuations of one (or a few) sites close to the boundary, while the contributions coming from the sites in the bulk of chains are self-averaging.
Concurrently, we have shown that the realization-dependent splitting probability and the realization-dependent mean first-passage time to a target placed at distance $N$ away from the starting point are self-averaging in the limit $N\to\infty$. Clearly, this result just indicates the trend in the behavior of fluctuations that they become progressively less important the farther is the target away from the starting point. However, for most of practical applications $N$ is essentially finite and our numerical analysis shows that for such $N$ fluctuations are quite significant.
@@ Acknowledgments
The authors wish to thank E. Agliari, T. Akimoto, D. S. Dean, E. Barkai and L. A. Pastur for helpful discussions. This project has been supported by funding from the 2021 first FIS (Fondo Italiano per la Scienza) funding scheme (FIS783—SMaC—Statistical Mechanics and Complexity University and Research) from Italian MUR (Ministry of University and Research).
@@ Appendix A Realization-dependent probability current
For the Model I the rates are defined in Eqs. (LABEL:rates), which gives
@@
from which we find the first expression in Eq. (21).
@@
Appendix B Realization-dependent splitting probability
In this Appendix we derive an explicit expression for the splitting probability $E_{-}$. We consider a continuous-time random walk on a finite one-dimensional lattice with sites $x=0,\,1,\,\dots,\,N$, with the transition rates $W_{x,\,x+1}$ and $W_{x,\,x-1}$ for jumps from the site $x$ to the neighboring sites x1x-1 and $x+1$. Let $E_{-}(x_{0})$ denote the probability that, starting from site $x_{0}$, the random walk visits the left boundary at $x=0$ before reaching the right boundary at $x=N$.
As shown in [77] the splitting probability $E_{-}(x)$ (here, the walk is assumed to start at arbitrary point $x$, excluding $x=0$ and $x=N$) satisfies the backward master equation of the form
$W_{x,x+1}\big[E_{-}(x+1)-E_{-}(x)\big]+W_{x,x-1}\big[E_{-}(x-1)-E_{-}(x)\big]=0,\quad x=1,\dots,N-1\,$, (100)
which is to be solved subject to trivial boundary conditions :
$E_{-}(x=0)=1,\qquad E_{-}(x=N)=0.$ (101)
Equation (100) can be conveniently rewritten in terms of finite differences $\Delta E_{-}(x)=E_{-}(x+1)-E_{-}(x)$,
$\Delta E_{-}(x)=\frac{W_{x,\,x-1}}{W_{x,\,x+1}}\,\Delta E_{-}(x-1),\qquad 1\leq x<N\,.$ (102)
and solved recursively to give
$\Delta E_{-}(x)=\Delta E_{-}(0)\prod_{k=1}^{x}\frac{W_{k,\,k-1}}{W_{k,\,k+1}}\,.$ (103)
The above expression yields the following explicit expression in terms of $E_{-}(x)$:
$E_{-}(x)=1+\Delta E_{-}(0)\Sigma_{x}\,$, (104)
in which $\Delta E_{-}(0)$ is as yet an undefined parameter and the sum $\Sigma_{x}$ has been introduced in A,
$\Sigma_{x}$ $=\displaystyle 1+\sum_{j=1}^{x-1}\prod_{k=1}^{j}\frac{W_{k,\,k-1}}{W_{k,\,k+1}}.$ (105)
Lastly, the boundary condition $E_{-}(N)=0$ ensures that
$\Delta E_{-}(0)=-\frac{1}{\Sigma_{N}},$
and hence, the general solution for $x=x_{0}$ reads
$E_{-}(x_{0})=1-\frac{\displaystyle\Sigma_{x_{0}}}{\displaystyle\Sigma_{N}}=\frac{\displaystyle\sum_{j=x_{0}}^{N-1}\prod_{k=1}^{j}\frac{W_{k,\,k-1}}{W_{k,\,k+1}}}{1+\displaystyle\sum_{j=1}^{N-1}\prod_{k=1}^{j}\frac{W_{k,\,k-1}}{W_{k,\,k+1}}}.$ (106)
We note next that the above expression (106) can be cast into a compact form in Eq. (53), where $\tau_{x_{0}}^{-}$ and $\tau_{N-x_{0}}^{+}$ represent, respectively, the resistances of the intervals $(0,x_{0}-1)$ and $(x_{0}+1,N)$ from the left and from the right of the starting point $x_{0}$:
$\begin{split}\displaystyle\tau_{x_{0}}^{-}&=\frac{1}{W_{x_{0},\,x_{0}-1}}\left(1+\frac{W_{x_{0}-1,\,x_{0}}}{W_{x_{0}-1,\,x_{0}-2}}+\dots+\frac{W_{x_{0}-1,\,x_{0}}\cdots W_{1,\,2}}{W_{x_{0}-1,\,x_{0}-2}\cdots W_{1,\,0}}\right),\\ \displaystyle\tau_{N-x_{0}}^{+}&=\frac{1}{W_{x_{0},\,x_{0}+1}}\left(1+\frac{W_{x_{0}+1,\,x_{0}}}{W_{x_{0}+1,\,x_{0}+2}}+\dots+\frac{W_{x_{0}+1,\,x_{0}}\cdots W_{N-1,\,N-2}}{W_{x_{0}+1,\,x_{0}+2}\cdots W_{N-1,\,N}}\right).\end{split}$ (107)
This is done by simply noticing that Eq. (106) can be formally rewritten as
[…]
and then factoring out the common product $W_{1,\,0}\cdots W_{x_{0},\,x_{0}-1}/W_{1,\,2}\cdots W_{x_{0},\,x_{0}+1}$ from both numerator and denominator.
Consider next the model-specific forms of the splitting probability $E_{-}(x_{0})$.
Model I. Using the general expressions derived in A, we have that the resistances of the intervals from the left and from the right from the starting site $x_{0}$ attain the following forms:
[…]
where $\phi_{x}$ is defined in the main text. This gives for the Model I the following expression
$E_{-}^{(I)}(x_{0})=\frac{\phi_{x_{0}}\phi_{x_{0}+1}+\phi_{x_{0}+1}\phi_{x_{0}+2}+\dots+\phi_{N-1}\phi_{N}}{\phi_{0}\phi_{1}+\dots+\phi_{N-1}\phi_{N}}=\frac{H_{x_{0},\,N}}{H_{0,N}}\,.$ (110)
Model II. In this case the resistances admit the following form
$\begin{split}\tau_{x_{0}}^{-}&=\frac{\phi_{x_{0}+1}+\phi_{x_{0}-1}}{\phi_{x_{0}+1}\phi_{x_{0}}\phi_{x_{0}-1}}\left(\phi_{x_{0}}\phi_{x_{0}-1}+\phi_{x_{0}-1}\phi_{x_{0}-2}+\dots+\phi_{1}\phi_{0}\right),\\[5.69054pt] \tau_{N-x_{0}}^{+}&=\frac{\phi_{x_{0}-1}+\phi_{x_{0}+1}}{\phi_{x_{0}-1}\phi_{x_{0}}\phi_{x_{0}+1}}\left(\phi_{x_{0}}\phi_{x_{0}+1}+\phi_{x_{0}+1}\phi_{x_{0}+2}+\dots+\phi_{N-1}\phi_{N}\right).\end{split}$ (111)
Noticing that the multipliers $(\phi_{x_{0}-1}+\phi_{x_{0}+1})/\phi_{x_{0}-1}\phi_{x_{0}}\phi_{x_{0}+1}$ cancel each other, we find that also for the Model II the splitting probability $E_{-}^{(II)}(x_{0})$ obeys Eq. (110).
Model III. For the random trap model the transition rates are symmetric and hence,
$\tau_{x_{0}}^{-}=\frac{x_{0}}{\phi_{x_{0}}^{2}}\,,\quad\tau_{N-x_{0}}^{+}=\frac{N-x_{0}}{\phi_{x_{0}}^{2}}.$ (112)
Consequently, we find that $E_{-}^{(III)}(x_{0})$ is independent of disorder (i.e., it is not a fluctuating property) and obeys
$E_{-}^{(III)}(x_{0})=1-\frac{x_{0}}{N}\,.$ (113)
Appendix C Realization-dependent diffusion coefficient
@@
Derrida’s result. For completeness, we present below the Derrida’s general result for the realization-dependent $D_{N}$ for random walks evolving in a continuous time on a finite periodic chain with $N$ sites with arbitrary fixed transition rates $W_{x,x\pm 1}$. As shown by Derrida [4], the realization-dependent diffusion coefficient in such a setting has the following form
@@
Model I. For the Model I the transition rates are defined in Eq. (LABEL:rates). Using this definition, we find
@@
where $H_{0,N}$ is given in Eq. (23). Further on, from Eq. (119) we have that
@@
Model II. In this dynamical scenario with randomized stepping-times, the transition rates $W_{x,x\pm 1}=p_{x,x\pm 1}/\delta t$, where $p_{x,x\pm 1}$ are the transition probabilities of an auxiliary random process that obey Eqs. (11) and (13). Firstly, we notice that
@@
Model III. In the Gaussian random trap model the transition rates obey Eq. (20). It is straightforward to verify that the product